OpenStudy (anonymous):

Solve a^2 - 81 = 0 by factoring. Select all possible answers. -9 0 9 No solution Infinite solutions

OpenStudy (anonymous):

A and C xD

OpenStudy (anonymous):

use a^2-b^2=(a+b)(a-b)

OpenStudy (haseeb96):

9 is the correct answer as u know 81 is the square of 9 so when u takeout the square root of 81 it will equal to 9

OpenStudy (anonymous):

Itsa Perfect square, you will have (a+9)(a-9). Then set each equal to zero.

OpenStudy (nurali):

(a+9)(a-9)=0

OpenStudy (anonymous):

you will get a^2-9^2=(a+9)(a-9) so a=9,-9

OpenStudy (nurali):

a=9 or a=-9

OpenStudy (anonymous):

So A and C.

OpenStudy (anonymous):

Thank you.

OpenStudy (haseeb96):